# arXiv:2602.14013v2[math.CV] 01 Mar 2026

### v2

@@ Abstract
The paper is devoted to the study of the global behavior of mappings. We consider families mappings with different definition domains which satisfy moduli inequalities of Poletskii type. Under some additional assumptions we have proved that such families are uniformly equicontinuous.
@@ 1 Introduction
This paper is devoted to the study of mappings satisfying upper distortion estimates of the modulus of families of paths. We consider the case when definition domains of mappings may be different. The main subject of this study is distortion estimates for the distance under mappings. For fixed domains, this problem has been repeatedly studied by many authors, including both homeomorphisms and mappings with a branching. In particular, we could cite papers on this matter, see, e.g., [Cr1][Cr2], [MRSY], [M1][M3], [RSY1][RSY2], [RS], [Sev1] and [SevSkv]. We note that we draw heavily on the ideas of Suvorov [Suv], who considered sequences of domains and the theory of prime ends for them. At the same time, the present paper is not, generally speaking, concerned with the theory of prime ends.
@@
holds for all $0<r_{1}<r_{2}<d_{0}:=\sup\limits_{x\in D}|x-x_{0}|$ and any Lebesgue measurable function $\eta:(r_{1},r_{2})\rightarrow[0,\infty]$ such that
@@
The inequalities of the form (1.3) were established for many well-known classes of mappings, see e.g. [MRV, Theorem 7.1], [Va, Definition 13.1]), [MRSY, Theorems 8.1, 8.5] and [Sal, Corollary 2].
@@
Let $I$ be a fixed set of indices and let $D_{i},$ $i\in I,$ be some number of domains. Following [NP, Sect. 2.4], we say that a family of domains $\{D_{i}\}_{i\in I}$ is equi-uniform with respect to $p$-modulus if for any $r>0$ there exists a number $\delta>0$ such that the inequality
$M_{p}(\Gamma(F^{\,*},F,D_{i}))\geqslant\delta$ (1.5)
holds for any $i\in I$ and any continua $F,F^{*}\subset D_{i}$ such that $h(F)\geqslant r$ and $h(F^{\,*})\geqslant r.$ Set
@@
Given $p\geqslant 1,$ a number $\delta>0,$ a fixed domain $D_{0}\subset{\mathbb{R}}^{n},$ $n\geqslant 2,$ a sequence of domains $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ the kernel of which is $D_{0},$ a non-degenerate continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{F}_{Q,A,p,\delta}(D_{0},\mathfrak{D})$ some a family of ring $Q$-homeomorphisms $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ for all $0<r_{1}<r_{2}<\infty$ such that $h(f_{m}(A))\geqslant\delta$ and $h(\overline{{\mathbb{R}}^{n}}\setminus f_{m}(D_{m}))\geqslant\delta.$
@@
Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$ In addition, assume that
5) $p=n$ and there is $r>0$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m}).$
Then $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ and some $\delta_{1}>0.$ Moreover, if $A_{0}$ is some another compactum in $D_{0},$ then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
@@
Given $p\geqslant 1,$ a domain $D_{0}\subset{\mathbb{R}}^{n},$ a set $E\subset\overline{{\mathbb{R}}^{n}}$ and a number $\delta>0$, a fixed domain $D_{0}\subset{\mathbb{R}}^{n},$ $n\geqslant 2,$ a sequence of domains $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ the kernel of which is $D_{0}$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$ we denote by $\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D})$ the family of all open discrete closed ring mappings $f:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}\setminus E$ satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ such that the following condition holds: for any domain $D^{\,\prime}_{m}:=f_{m}(D_{m})$ there exists a continuum $K_{m}\subset D^{\,\prime}_{m}$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ The following statement holds.
@@
Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$ Besides that, if the condition 5) of Theorem 1 holds, then for any continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ such that $h(f_{m}(A))\geqslant\delta$ for any $m=1,2,\ldots,$ there exists $\delta_{1}>0$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$
@@
The next definitions due to Caratheodory [Car]; cf. [KR]. Let $\omega$ be an open set in ${\mathbb{R}}^{k}$, $k=1,\ldots,n-1$. A continuous mapping $\sigma\colon\omega\rightarrow{\mathbb{R}}^{n}$ is called a $k$-dimensional surface in ${\mathbb{R}}^{n}$. A surface is an arbitrary (n1)(n-1)-dimensional surface $\sigma$ in ${\mathbb{R}}^{n}.$ A surface $\sigma$ is called a Jordan surface, if $\sigma(x)\neq\sigma(y)$ for $x\neq y$. In the following, we will use $\sigma$ instead of $\sigma(\omega)\subset{\mathbb{R}}^{n},$ $\overline{\sigma}$ instead of $\overline{\sigma(\omega)}$ and $\partial\sigma$ instead of $\overline{\sigma(\omega)}\setminus\sigma(\omega).$ A Jordan surface $\sigma\colon\omega\rightarrow D$ is called a cut of $D$, if $\sigma$ separates $D,$ that is $D\setminus\sigma$ has more than one component, $\partial\sigma\cap D=\varnothing$ and $\partial\sigma\cap\partial D\neq\varnothing$.
A sequence of cuts $\sigma_{1},\sigma_{2},\ldots,\sigma_{m},\ldots$ in $D$ is called a chain, if:
(i) the set $\sigma_{m+1}$ is contained in exactly one component $d_{m}$ of the set $D\setminus\sigma_{m},$ wherein $\sigma_{m-1}\subset D\setminus(\sigma_{m}\cup d_{m})$; (ii) $\bigcap\limits_{m=1}^{\infty}\,d_{m}=\varnothing.$
Two chains of cuts $\{\sigma_{m}\}$ and $\{\sigma_{k}^{\,\prime}\}$ are called equivalent, if for each $m=1,2,\ldots$ the domain $d_{m}$ contains all the domains $d_{k}^{\,\prime}$, except for a finite number, and for each $k=1,2,\ldots$ the domain $d_{k}^{\,\prime}$ also contains all domains $d_{m},$ except for a finite number.
The end of the domain $D$ is the class of equivalent chains of cuts in $D$. Let $K$ be the end of $D$ in ${\mathbb{R}}^{n}$, then the set $I(K)=\bigcap\limits_{m=1}\limits^{\infty}\overline{d_{m}}$ is called the impression of $K$. Following [Na], we say that the end $K$ is a prime end, if $K$ contains a chain of cuts $\{\sigma_{m}\}$ such that
$\lim\limits_{m\rightarrow\infty}M(\Gamma(C,\sigma_{m},D))=0$
for some continuum $C$ in $D.$ The following notation is used: the set of prime ends corresponding to the domain $D,$ is denoted by $E_{D},$ and the completion of the domain $D$ by its prime ends is denoted $\overline{D}_{P}.$
Consider the following definition, which goes back to Näkki [Na], cf. [KR]. The boundary of a domain $D$ in ${\mathbb{R}}^{n}$ is said to be locally quasiconformal if every $x_{0}\in\partial D$ has a neighborhood $U$ that admits a quasiconformal mapping $\varphi$ onto the unit ball ${\mathbb{B}}^{n}\subset{\mathbb{R}}^{n}$ such that $\varphi(\partial D\cap U)$ is the intersection of ${\mathbb{B}}^{n}$ and a coordinate hyperplane. The sequence of cuts $\sigma_{m},$ $m=1,2,\ldots,$ is called quasiconformally regular, if $\overline{\sigma_{m}}\cap\overline{\sigma_{m+1}}=\varnothing$ for $m\in{\mathbb{N}}$ and, in addition, $d(\sigma_{m})\rightarrow 0$ as $m\rightarrow\infty.$ If the end $K$ contains at least one quasiconformally regular chain, then $K$ will be called quasiconformally regular. We say that a bounded domain $D$ in ${\mathbb{R}}^{n}$ is quasiconformally regular, if $D$ can be quasiconformally mapped to a domain with a locally quasiconformal boundary whose closure is a compact in ${\mathbb{R}}^{n},$ and, besides that, every prime end in $D$ is quasiconformally regular. Note that space $\overline{D}_{P}=D\cup E_{D}$ is metric, which can be demonstrated as follows. If $g:D_{0}\rightarrow D$ is a quasiconformal mapping of a domain $D_{0}$ with a locally quasiconformal boundary onto some domain $D,$ then for $x,y\in\overline{D}_{P}$ we put:
@@
where the element $g^{\,-1}(x),$ $x\in E_{D},$ is to be understood as some (single) boundary point of $D_{0}.$ The specified boundary point is unique and well-defined by [IS, Theorem 2.1, Remark 2.1], cf. [Na, Theorem 4.1]. It is easy to verify that $\rho$ in (1.7) is a metric on $\overline{D}_{P}.$ If $g_{*}$ is another quasiconformal mapping of a domain $D_{*}$ with locally quasiconformal boundary onto $D$, then the corresponding metric $\rho_{*}(p_{1},p_{2})=|{\widetilde{g_{*}}}^{-1}(p_{1})-{\widetilde{g_{*}}}^{-1}(p_{2})|$ generates the same convergence and, consequently, the same topology in $\overline{D}_{P}$ as $\rho_{0}$ because $g_{0}\circ g_{*}^{-1}$ is a quasiconformal mapping of $D_{*}$ and $D_{0}$, which extends, by Theorem 4.1 in [Na], to a homeomorphism between $\overline{D_{*}}$ and $\overline{D_{0}}$. In the sequel, this topology in $\overline{D}_{P}$ will be called the topology of prime ends; the continuity of mappings $F\colon\overline{D}_{P}\rightarrow\overline{D^{\,\prime}}_{P}$ will be understood relative to this topology.
We say that the boundary of $D$ is weakly flat at a point $x_{0}\in\partial D$ if, for every number $P>0$ and every neighborhood $U$ of the point $x_{0},$ there is a neighborhood $V\subset U$ such that $M_{\alpha}(\Gamma(E,F,D))\geqslant P$ for all continua $E$ and $F$ in $D$ intersecting $\partial U$ and $\partial V.$ We say that the boundary $\partial D$ is weakly flat if the corresponding property holds at every point of the boundary.
@@
Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D}$ in terms of prime ends, i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $\rho(x,y)<\delta,$ $m\in{\mathbb{N}}$ and $\rho$ is one of the metric in $\overline{D_{0}}_{P}.$
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}$ with respect $\rho.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}_{P}\rightarrow\overline{{\mathbb{R}}^{n}}.$ In addition, assume that
5) $p=n$ and there is $r>0$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m}).$
Then $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ and some $\delta_{1}>0.$ Moreover, if $A_{0}$ is some another compactum in $D_{0},$ then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
@@
Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D}$ with respect to prime ends, i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $\rho(x,y)<\delta,$ $m\in{\mathbb{N}}$ and $\rho$ is one of the metric in $\overline{D_{0}}_{P}.$
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$ Besides that, if the condition 5) of Theorem 1 holds, then for any continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ such that $h(f_{m}(A))\geqslant\delta$ for any $m=1,2,\ldots,$ there exists $\delta_{1}>0$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$

### v3

# ON DISTORTION ESTIMATES UNDER MAPPINGS OF FAMILIES OF DOMAINS

@@ Abstract
We consider families of mappings with moduli inequalities, having different definition domains. Under some additional assumptions we have proved that such families are uniformly equicontinuous. We have considered four main cases: when mappings are homeomorphisms and corresponding domains have simple geometry; when similar mappings have branch points; when domains with complex geometry are considered, but mappings still are homeomorphisms; and when similar mappings have branch points. Sequences of domains are generally assumed to converge to a kernel, and the characteristics of the mappings must satisfy certain conditions on their growth. In some of the four cases mentioned above, we also described properties of the limit mapping. We also obtained the correspondence of the boundary points of the kernel to the boundary points, and the inner points to the inner points.
@@ 1 Introduction
This paper is devoted to the study of mappings satisfying upper distortion estimates of the modulus of families of paths. We consider the case when definition domains of mappings may be different. The main subject of this study is distortion estimates for the distance under mappings. For fixed domains, this problem has been studied by many authors, including both homeomorphisms and mappings with a branching. In particular, we could mentioned papers on this matter, see, e.g., [Cr1][Cr2], [MRSY], [M], [RS1] and [SevSkv]. Considering families of mappings of variable domains, we use some ideas from [Suv].
@@
holds for all $0<r_{1}<r_{2}<\infty$ and any Lebesgue measurable function $\eta:(r_{1},r_{2})\rightarrow[0,\infty]$ such that
@@
The inequalities of the form (1.3) were established for many well-known classes of mappings, see e.g. [Va, Definition 13.1]), [MRSY, Theorems 8.1, 8.5] and [Sal, Corollary 2].
@@
Let $I$ be a fixed set of indices and let $D_{i},$ $i\in I,$ be some number of domains. Following [NP, Sect. 2.4], we say that a family of domains $\{D_{i}\}_{i\in I}$ is equi-uniform with respect to $p$-modulus if for any $r>0$ there exists a number $\delta>0$ such that the inequality $M_{p}(\Gamma(F^{\,*},F,D_{i}))\geqslant\delta$ holds for any $i\in I$ and any continua $F,F^{*}\subset D_{i}$ such that $h(F)\geqslant r$ and $h(F^{\,*})\geqslant r.$
Remark 1.1. Due to [MRSY, Section 3] we say that a boundary $D$ is called strongly accessible with respect to $p$-modulus at $x_{0}\in\partial D,$ if for any neighborhood $U$ of the point $x_{0}\in\partial D$ there is a neighborhood $V\subset U$ of this point, a compactum $F\subset D$ and a number $\delta>0$ such that $M_{p}(\Gamma(E,F,D))\geqslant\delta$ for any continua $E\subset D$ such that $E\cap\partial U\neq\varnothing\neq E\cap\partial V.$ The boundary of a domain $D$ is called strongly accessible with respect to $p$-modulus, if this is true for any $x_{0}\in\partial D.$ When $p=n,$ prefix ‘‘relative to $p$-modulus’’ is omitted. Uniform domains have strongly accessible boundaries, see e.g. [SevSkv, Remark 1]. The inverse statement is also true at least for $p=n,$ see [Sev2, Lemma 4.2].
Remark 1.2. Observe that, if $n=2,$ then every domain $D$ in ${\mathbb{R}}^{2}$ which has a finite number of boundary components is uniform if and only if $D$ is finitely connected on the boundary, see Corollary 6.8 in [Na2].
Set
@@
Given $p\geqslant 1,$ a number $\delta>0,$ a fixed domain $D_{0}\subset{\mathbb{R}}^{n},$ $n\geqslant 2,$ a sequence of domains $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ the kernel of which is $D_{0},$ a non-degenerate continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{F}_{Q,A,p,\delta}(D_{0},\mathfrak{D})$ some a family of ring $Q$-homeomorphisms $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ for all $0<r_{1}<r_{2}<\infty$ such that $h(f_{m}(A))\geqslant\delta$ and $h(\overline{{\mathbb{R}}^{n}}\setminus f_{m}(D_{m}))\geqslant\delta.$ The concept of finite mean oscillation (FMO) used below can be found, for example, in [MRSY] or [Sev1]. The following statement is true.
@@
I. Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$
II. Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}$ and $f$ is a homeomorphism in $D_{0}.$ If $A_{0}$ is some compactum in $D_{0},$ then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
III. If $x_{m}\in D_{m},$ $m=1,2,\ldots,$ $f_{m}$ converges to $f$ locally uniformly in $D_{0}$ and $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty,$ then $f_{m}(x_{m})\rightarrow f(x_{0}).$ If $x_{0}\in\partial D_{0},$ then $f(x_{0})\in\partial f(D_{0}).$
@@
Given $p\geqslant 1,$ a domain $D_{0}\subset{\mathbb{R}}^{n},$ a set $E\subset\overline{{\mathbb{R}}^{n}}$ and a number $\delta>0$, a fixed domain $D_{0}\subset{\mathbb{R}}^{n},$ $n\geqslant 2,$ a sequence of domains $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ the kernel of which is $D_{0}$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$ we denote by $\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D})$ the family of all open discrete closed mappings $f:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}\setminus E$ satisfying (1.3)–(1.4) for any $0<r_{1}<r_{2}<\infty$ at any point $x_{0}\in\overline{D_{0}}$ such that the following condition holds: for any domain $D^{\,\prime}_{m}:=f_{m}(D_{m})$ there exists a continuum $K_{m}\subset D^{\,\prime}_{m}$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ The following statement holds.
@@
I. Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$ Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$ If $x_{m}\in D_{m},$ $m=1,2,\ldots,$ $f_{m}$ converges to $f$ locally uniformly in $D_{0}$ and $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty,$ then $f_{m}(x_{m})\rightarrow f(x_{0}).$
II. In addition, assume that
5) $p=n$ and there is $r>0,$ does not depending on $m\in{\mathbb{N}},$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m}).$ If $A_{0}$ is some compactum in $D_{0},$ then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$ The mapping $f$ is boundary preserving: if $x_{0}\in\partial D_{0},$ then $f(x_{0})\in\partial f(D_{0}).$ If $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}\cap\bigcap\limits_{m=1}^{\infty}D_{m},$ then there is $\varepsilon_{1}>0$ does not depending on $m\in{\mathbb{N}}$ such that
$B_{h}(f_{m}(x_{0}),\varepsilon_{1})\subset f_{m}(D_{m})\,,\quad m=1,2,\ldots\,$, (1.6)
where $B_{h}(y_{0},r_{0})=\{y\in\overline{{\mathbb{R}}^{n}}:h(y,y_{0})<r_{0}\}.$
@@ II. In addition, assume that
Definitions related to prime ends of domains can be found in [Sev1] and are therefore omitted. Consider the following definition, which goes back to Näkki [Na1], cf. [KR]. The boundary of a domain $D$ in ${\mathbb{R}}^{n}$ is said to be locally quasiconformal if every $x_{0}\in\partial D$ has a neighborhood $U$ that admits a quasiconformal mapping $\varphi$ onto the unit ball ${\mathbb{B}}^{n}\subset{\mathbb{R}}^{n}$ such that $\varphi(\partial D\cap U)$ is the intersection of ${\mathbb{B}}^{n}$ and a coordinate hyperplane. The sequence of cuts $\sigma_{m},$ $m=1,2,\ldots,$ is called quasiconformally regular, if $\overline{\sigma_{m}}\cap\overline{\sigma_{m+1}}=\varnothing$ for $m\in{\mathbb{N}}$ and, in addition, $d(\sigma_{m})\rightarrow 0$ as $m\rightarrow\infty.$ If the end $K$ contains at least one quasiconformally regular chain, then $K$ will be called quasiconformally regular. We say that a bounded domain $D$ in ${\mathbb{R}}^{n}$ is quasiconformally regular, if $D$ can be quasiconformally mapped to a domain with a locally quasiconformal boundary whose closure is a compact in ${\mathbb{R}}^{n},$ and, besides that, every prime end in $D$ is quasiconformally regular. Note that space $\overline{D}_{P}=D\cup E_{D}$ is metric, which can be demonstrated as follows. If $g:D_{0}\rightarrow D$ is a quasiconformal mapping of a domain $D_{0}$ with a locally quasiconformal boundary onto some domain $D,$ then for $x,y\in\overline{D}_{P}$ we put:
@@
where the element $g^{\,-1}(x),$ $x\in E_{D},$ is to be understood as some (single) boundary point of $D_{0}.$ The specified boundary point is unique and well-defined by [IS, Theorem 2.1, Remark 2.1], cf. [Na1, Theorem 4.1]. It is easy to verify that $\rho$ in (1.7) is a metric on $\overline{D}_{P}.$ If $g_{*}$ is another quasiconformal mapping of a domain $D_{*}$ with locally quasiconformal boundary onto $D$, then the corresponding metric $\rho_{*}(p_{1},p_{2})=|{\widetilde{g_{*}}}^{-1}(p_{1})-{\widetilde{g_{*}}}^{-1}(p_{2})|$ generates the same convergence and, consequently, the same topology in $\overline{D}_{P}$ as $\rho_{0}$ because $g_{0}\circ g_{*}^{-1}$ is a quasiconformal mapping of $D_{*}$ and $D_{0}$, which extends, by Theorem 4.1 in [Na1], to a homeomorphism between $\overline{D_{*}}$ and $\overline{D_{0}}$. In the sequel, this topology in $\overline{D}_{P}$ will be called the topology of prime ends; the continuity of mappings $F\colon\overline{D}_{P}\rightarrow\overline{D^{\,\prime}}_{P}$ will be understood relative to this topology.
@@
I. Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D}$ in terms of prime ends, i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $\rho(x,y)<\delta,$ $m\in{\mathbb{N}}$ and $\rho$ is one of the metric in $\overline{D_{0}}_{P}$ defined in (1.7).
II. Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}$ with respect $\rho.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}_{P}\rightarrow\overline{{\mathbb{R}}^{n}}$ and $f$ is a homeomorphism in $D_{0}.$ If $A_{0}$ is some compactum in $D_{0},$ then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
III. If $x_{m}\in D_{m},$ $m=1,2,\ldots,$ $f_{m}$ converges to $f$ locally uniformly in $D_{0}$ and $x_{m}\rightarrow P_{0}$ as $m\rightarrow\infty,$ then $f_{m}(x_{m})\rightarrow f(P_{0}).$ If $P_{0}\in E_{D_{0}}:=\overline{D_{0}}_{P}\setminus D_{0},$ then $f(P_{0})\in\partial f(D_{0}).$
@@
I. Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D}$ with respect to prime ends, i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $\rho(x,y)<\delta,$ $m\in{\mathbb{N}}$ and $\rho$ is one of the metric in $\overline{D_{0}}_{P}.$ Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}_{P}\rightarrow\overline{{\mathbb{R}}^{n}}.$ If $x_{m}\in D_{m},$ $m=1,2,\ldots,$ $f_{m}$ converges to $f$ locally uniformly in $D_{0}$ and $x_{m}\rightarrow P_{0}$ as $m\rightarrow\infty,$ then $f_{m}(x_{m})\rightarrow f(P_{0}).$
II. Besides that, if the condition 5) of Theorem 1 holds, then for any compactum $A_{0}\subset D_{0},$ $A_{0}\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ there exists $\delta_{1}>0$ such that $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$ The mapping $f$ is boundary preserving: if $P_{0}\in E_{D_{0}},$ then $f(x_{0})\in\partial f(D_{0}).$ If $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}\cap\bigcap\limits_{m=1}^{\infty}D_{m},$ then there is $\varepsilon_{1}>0$ such that the relation (1.6) holds, where $B_{h}(y_{0},r_{0})=\{y\in\overline{{\mathbb{R}}^{n}}:h(y,y_{0})<r_{0}\}.$